### v1

@@ 6 Proofs
Since this theorem states the minimax risk of convergence, we will first discuss the maximal risk of an estimator and then proceed with the matching lower bound.
Argument for the upper bound
Consider the conventional learning estimator $\widehat{f}^{[t]}$ which is the output of algorithm $\mathcal{A}_{\mathrm{CL}}(b_{t},d_{t},M_{t})$ with the specifications in Theorem 2.2. We would argue that it attains the optimal rate of convergence in Theorem 2.1. The first step is to conceive a “good” piecewise polynomial approximation $\bar{f}^{[t]}$ of target mean function $f^{[t]}$. Let us partition $[0,1]$ into $b_{t}^{-1}$-many intervals of length $b_{t}$. These are denoted by $I_{r}^{[t]}$ $(r=1,\ldots,b_{t}^{-1})$ from left to right. Since we have assumed the smoothness, $f^{[t]}\in\mathcal{H}_{\alpha_{m}}(L_{m},M_{m})$, it is straightforward to take $\bar{f}^{[t]}:[0,1]\to\mathbb{R}$ by
$\bar{f}^{[t]}(x):=\sum_{r=1}^{b_{t}^{-1}}\sum_{q=0}^{\omega(\alpha_{m})}\mathbb{1}(x\in I_{r}^{[t]})\frac{\bigl{(}x-(r-1)b_{t}\bigr{)}^{q}}{q!}\frac{d^{q}f^{[t]}}{dx^{q}}\bigl{(}x-(r-1)b_{t}\bigr{)},\quad(0\leq x\leq 1).$
By Taylor’s theorem, it would be immediate to see
$\lVert\bar{f}^{[t]}-f^{[t]}\rVert_{\mathcal{L}^{\infty}}\leq\lambda_{t}\qquad\text{where}~{}\lambda_{t}:=\frac{L_{m}}{\bigl{(}\omega(\alpha_{m})\bigr{)}!}b_{t}^{\alpha_{m}}.$
Our approach is to establish an upper bound for the right-hand side of the following:
$\lvert\widehat{f}^{[t]}(x)-f^{[t]}(x)\rvert\leq\lambda_{t}+\lvert\widehat{f}^{[t]}(x)-\bar{f}^{[t]}(x)\rvert,\qquad\text{for every}~{}0\leq x\leq 1.$
Let us denote by $\check{f}^{[t]}$ the local polynomial fit of degree $d_{t}$ right before thresholding in the algorithm $\mathcal{A}(\mathcal{D}^{[t]},b_{t},d_{t},M_{t})$. Since the estimator $\widehat{f}^{[t]}$ is just a thresholded version of $\check{f}^{[t]}$, both $\widehat{f}^{[t]}$ and $\bar{f}^{[t]}$ have essentially the same forms throughout the interval $I_{r}^{[t]}$ $(r=1,\ldots,b_{t}^{-1})$. Now that the difference $\widehat{f}^{[t]}(x)-\bar{f}^{[t]}(x)$ is of main interest, we shall focus on the specific interval $I_{1}^{[t]}=[0,b_{t})$ without loss of generality. The results on other intervals are obvious by symmetry.
We shall denote by $\bigl{\{}(\mathbb{T}_{1}^{[t]},\mathbb{Y}_{1}^{[t]}),\ldots,(\mathbb{T}_{n_{t}}^{[t]},\mathbb{Y}_{n_{t}}^{[t]})\bigr{\}}$ the random collection utilized for deriving the intermediate estimator $\check{f}^{[t]}$ with polynomial regression on the interval $I_{1}^{[t]}$. The design points $\mathbb{T}_{1}^{[t]},\ldots,\mathbb{T}_{n_{t}}^{[t]}$ must be independent and supported on some discrete set $\mathcal{T}$. Notice that $\check{f}^{[t]}$ is a well-defined polynomial fit on $I_{1}^{[t]}$ if the number of target subjects is large enough as $n_{t}\geq d_{t}+1$ and there are at least $d_{t}+1$ distinct design points among $\mathbb{T}_{1}^{[t]},\ldots,\mathbb{T}_{n_{t}}^{[t]}$. By assuming $b_{t}\geq 2C_{t}(d_{t}+1)/m_{t}$, such an event, denoted by $A_{1}$, occurs with probability at least $1-2^{-n_{t}}$.
Let us define a vector $u_{t}(x):=(1,x/b_{t},\ldots,(x/b_{t})^{d_{t}})^{\top}\in\mathbb{R}^{d_{t}+1}$ for each $x\in I_{1}^{[t]}$. Under the event $A_{1}$, we have
$\check{f}^{[t]}(x)$ $=u_{t}(x)^{\top}\check{\beta}_{t}$ $\text{where}~{}\begin{dcases}W_{t}:=\bigl{(}u_{t}(\mathbb{T}_{1}^{[t]}),\ldots,u_{t}(\mathbb{T}_{n_{t}}^{[t]})\bigr{)}^{\top}\in\mathbb{R}^{n_{t}\times(d_{t}+1)},\\ y_{t}:=\bigl{(}\mathbb{Y}_{1}^{[t]},\ldots,\mathbb{Y}_{n_{t}}^{[t]}\bigr{)}^{\top}\in\mathbb{R}^{n_{t}},\\ \check{\beta}_{t}:=(W_{t}^{\top}W_{t})^{-1}W_{t}^{\top}y_{t}\in\mathbb{R}^{d_{t}+1},\\ \end{dcases}$
$\bar{f}^{[t]}(x)$ $=u_{t}(x)^{\top}\bar{\beta}_{t}$ for some appropriate $\text{for some appropriate vector}~{}\bar{\beta}_{t}\in\mathbb{R}^{d_{t}+1}.$
We introduce a useful lemma regarding the random matrix $(W_{t}^{\top}W_{t})^{-1}$, with a detailed proof available in the Supplementary material [5]. Let us denote by $A_{2}$ the event inside the probability on the left-hand side of Equation (6.1).
Lemma 6.1.
As long as $m_{t}b_{t}\geq 2C_{t}$ and $d_{t}\leq O(1)$ are true, there must be two generic constants $C_{1},C_{2}>0$ such that
$\mathbb{P}\Biggl{(}\biggl{\lVert}\biggl{(}\frac{W_{t}^{\top}W_{t}}{n_{t}}\biggr{)}^{-1}\biggr{\rVert}_{2}\leq C_{1}\Biggr{)}\geq 1-O(e^{-C_{2}\log^{2}n_{t}}).$
On the other hand, thanks to our model construction, the random vector $y_{t}$ is decomposed as $y_{t}=z_{t}+\varepsilon_{t}$ for random vectors $z_{t},\varepsilon_{t}\in\mathbb{R}^{n_{t}}$ such that
$z_{t}=\begin{bmatrix}f^{[t]}(\mathbb{T}_{1}^{[t]})\\ \vdots\\ f^{[t]}(\mathbb{T}_{n_{t}}^{[t]})\end{bmatrix},\quad\text{and}\quad\varepsilon_{t}=\begin{bmatrix}\mathbb{Y}_{1}^{[t]}-f^{[t]}(\mathbb{T}_{1}^{[t]})\\ \vdots\\ \mathbb{Y}_{n_{t}}^{[t]}-f^{[t]}(\mathbb{T}_{n_{t}}^{[t]})\end{bmatrix}.$
Due to our assumptions, $\lVert z_{t}\rVert_{\infty}\leq L_{m}$ and entries of $\varepsilon_{t}$ are independent sub-Gaussian random variables with variance proxy $2\tau^{2}$. The sub-Gaussian property yields the high-probability result as follows (e.g., Wainwright, [36]). Now $A_{3}$ denotes the high-probability event on the left-hand side of Equation (6.1).
$\mathbb{P}\bigl{(}\lVert\varepsilon_{t}\rVert_{\infty}\leq 2\sqrt{2}\tau\log n_{t}\bigr{)}\geq 1-2e^{-\log^{2}n_{t}}.$
From now on, we proceed with the discussion under the event $A_{1}\cap A_{2}\cap A_{3}$ whose probability should be at least $1-O(e^{-C_{3}\log^{2}n_{t}})$ for some generic constant $C_{3}>0$. If $n_{t}$ is large enough, we may easily notice that for any $x\in I_{1}^{[t]}$,
$\lvert\check{f}^{[t]}(x)\rvert$ $\leq\lVert u_{t}(x)\rVert_{2}\lVert(W_{t}^{\top}W_{t})^{-1}\rVert_{2}\lVert W_{t}^{\top}y_{t}\rVert_{2}$
$\leq C_{1}(d_{t}+1)n_{t}^{-1}\lVert W_{t}\rVert_{F}\lVert z_{t}+\varepsilon_{t}\rVert_{\infty}$
$\leq C_{1}(d_{t}+1)^{3/2}n_{t}^{-1/2}(L_{m}+2\sqrt{2}\tau\log n_{t})\leq\log n_{t}.$
This means our estimator $\widehat{f}^{[t]}$ is exactly the same as $\check{f}^{[t]}$ on the interval $I_{1}^{[t]}$. By Equation (6.1),
$\lvert\widehat{f}^{[t]}(x)-f^{[t]}(x)\rvert\leq\lambda_{t}+\lvert\check{f}^{[t]}(x)-\bar{f}^{[t]}(x)\rvert,\qquad\text{for all}~{}~{}x\in I_{1}^{[t]}.$
Next, let us focus on the second term of the right-hand side. The random vector $z_{t}$ is further decomposed as $z_{t}=W_{t}\bar{\beta}_{t}+\zeta_{t}$ for a random $\zeta_{t}\in\mathbb{R}^{n_{t}}$ such that
$\zeta_{t}=\begin{bmatrix}(f^{[t]}-\bar{f}^{[t]})(\mathbb{T}_{1}^{[t]})\\ \vdots\\ (f^{[t]}-\bar{f}^{[t]})(\mathbb{T}_{n_{t}}^{[t]})\end{bmatrix}.$
Considering the approximation result in Equation (6.1), we know $\lVert\zeta_{t}\rVert_{\infty}\leq\lambda_{t}$. On top of that, for every $x\in I_{1}^{[t]}$, it is trivial to see from $\bar{f}^{[t]}(x)=u_{t}(x)^{\top}\bar{\beta}_{t}$ to check that
$\check{f}^{[t]}(x)-\bar{f}^{[t]}(x)=u_{t}(x)^{\top}(W_{t}^{\top}W_{t})^{-1}W_{t}^{\top}(\zeta_{t}+\varepsilon_{t}).$
An upper bound for this quantity could be established as below: for any $x\in I_{1}^{[t]}$,
$\lvert\check{f}^{[t]}(x)-\bar{f}^{[t]}(x)\rvert$ $\leq\lVert u_{t}(x)\rVert_{2}\lVert(W_{t}^{\top}W_{t})^{-1}\rVert_{2}\lVert W_{t}^{\top}(\zeta_{t}+\varepsilon_{t})\rVert_{2}$
$\leq C_{1}(d_{t}+1)n_{t}^{-1}\lVert W_{t}\rVert_{F}\lVert\zeta_{t}+\varepsilon_{t}\rVert_{\infty}$
$\leq C_{1}(d_{t}+1)^{3/2}n_{t}^{-1/2}(\lambda_{t}+2\sqrt{2}\tau\log n_{t}).$
As far as the event $A_{1}\cap A_{2}\cap A_{3}$ occurs and both $m_{t}b_{t}\geq 2C_{t}$ and $\omega(\alpha_{m})\leq d_{t}\leq O(1)$ are true, we therefore have shown that
$\lVert\widehat{f}^{[t]}-f^{[t]}\rVert_{\mathcal{L}^{\infty}(I_{1}^{[t]})}\leq O\!\left(\lambda_{t}+\frac{\log n_{t}}{\sqrt{n_{t}}}\right).$
Recall that the event $A_{1}\cap A_{2}\cap A_{3}$ occurs with probability at least $1-O(e^{-C_{3}\log^{2}n_{t}})$ as well as the thresholding step implies $\lVert\widehat{f}^{[t]}-f^{[t]}\rVert_{\mathcal{L}^{\infty}(I_{1}^{[t]})}\leq O(\log n_{t})$. It is straightforward to obtain that
$\frac{1}{b_{t}}\mathbb{E}\lVert\widehat{f}^{[t]}-f^{[t]}\rVert_{\mathcal{L}^{2}(I_{1}^{[t]})}^{2}$
$~{}~{}\leq\mathbb{E}\Bigl{(}\lVert\widehat{f}^{[t]}-f^{[t]}\rVert_{\mathcal{L}^{\infty}(I_{1}^{[t]})}^{2}\mathbb{1}_{A_{1}\cap A_{2}\cap A_{3}}\Bigr{)}+\mathbb{E}\Bigl{(}\lVert\widehat{f}^{[t]}-f^{[t]}\rVert_{\mathcal{L}^{\infty}(I_{1}^{[t]})}^{2}\mathbb{1}_{A_{1}^{c}\cup A_{2}^{c}\cup A_{3}^{c}}\Bigr{)}$
$~{}~{}\leq O\!\left(\lambda_{t}+\frac{\log n_{t}}{\sqrt{n_{t}}}\right)^{\!2}+O(e^{-C_{3}\log^{2}n_{t}}\log^{2}n_{t})\leq O\!\left(\lambda_{t}^{2}+\frac{\log^{2}n_{t}}{n_{t}}\right).$
Now that our construction is symmetric throughout all other intervals $I_{q}^{[t]}$ $(q=2,\ldots,b_{t}^{-1})$,
$\sup_{\mathbb{P}\in\mathcal{P}}\mathbb{E}\lVert\widehat{f}^{[t]}-f^{[t]}\rVert_{\mathcal{L}^{2}}^{2}$ $\leq\sum_{r=1}^{b_{t}^{-1}}\sup_{\mathbb{P}\in\mathcal{P}}\mathbb{E}\lVert\widehat{f}^{[t]}-f^{[t]}\rVert_{\mathcal{L}^{2}(I_{q}^{[t]})}^{2}$
$\leq O\left(\lambda_{t}^{2}+\frac{\log^{2}n_{t}}{n_{t}}\right)\leq O\left(L_{m}^{2}b_{t}^{2\alpha_{m}}+\frac{\log^{2}n_{t}}{n_{t}}\right).$
The desired upper bound is now immediate when a bandwidth $b_{t}=\Theta(m_{t}^{-1})$ is selected to satisfy $m_{t}b_{t}\geq 2C_{t}(d_{t}+1)$, and a constant degree is chosen with $d_{t}\geq\omega(\alpha_{m})$.
Argument for the lower bound
Since there is no source sample $(n_{s}=0)$, we are enough to focus on the distribution of the target sample. Suppose that target random function $X^{[t]}_{1}$ is an unknown constant function. In this situation, the problem is fundamentally equivalent to estimating the scalar mean from $n_{t}$ number of independent and identically distributed observations under mean squared error. It is thus immediate to obtain the lower bound of the parametric rate:
$\sup_{\mathbb{P}\in\mathcal{P}}\mathbb{E}\lVert\widehat{f}^{[t]}-f^{[t]}\rVert^{2}_{\mathcal{L}^{2}}\gtrsim\frac{1}{n_{t}}.$
We are now enough to argue the other part:
$\sup_{\mathbb{P}\in\mathcal{P}}\mathbb{E}\lVert\widehat{f}^{[t]}-f^{[t]}\rVert^{2}_{\mathcal{L}^{2}}\gtrsim L_{m}^{2}m_{t}^{-2\alpha_{m}}.$
Let us consider any $C^{\infty}$ function $\phi:\mathbb{R}\to\mathbb{R}$ such that $\phi^{(q)}(x)=0$ for any $q=0,1,2,\ldots$ and $x\not\in(0,1)$. One of the common examples would be
$\phi(x)=\exp\left(-\frac{1}{t(1-t)}\right)\mathbb{1}(0<t<1)\qquad(t\in\mathbb{R}).$
For each $v=(v_{1},\ldots,v_{2m_{t}})\in\{0,1\}^{2m_{t}}$, we define a function $g_{v}:[0,1]\to\mathbb{R}$ by
$g_{v}(x):=\sum_{r=1}^{2m_{t}}CL_{m}v_{r}(2m_{t})^{-\alpha_{m}}\phi\bigl{(}2m_{t}x-(r-1)\bigr{)},$
where $C>0$ is a small enough generic constant such that $g_{v}\in\mathcal{H}_{\alpha_{m}}(L_{m},M_{m})$. We are ready to define sub-model $\mathcal{P}_{0}\subset\mathcal{P}$ which consists of a probability measure $\mathbb{P}_{v}$ with $v\in\{0,1\}^{2m_{t}}$. It could be any distribution as long as the target mean function $f^{[t]}$ is given as $g_{v}$. This sub-model $\mathcal{P}_{0}$ is Hamming separated in the sense that for any $v,w\in\{0,1\}^{2m_{t}}$,
$\lVert g_{v}-g_{w}\rVert_{\mathcal{L}^{2}}^{2}\geq C^{2}L_{m}^{2}(2m_{t})^{-2\alpha_{m}-1}\lVert\phi\rVert_{\mathcal{L}^{2}}^{2}\sum_{r=1}^{2m_{t}}\mathbb{1}(v_{r}\not=w_{r}).$
For each $r=1,\ldots,2m_{t}$, let ¯$\overline{\mathbb{P}}_{0,r}$ and ¯$\overline{\mathbb{P}}_{1,r}$ denote the mixture distributions of $\{\mathbb{P}_{v}:v_{r}=0\}$ and $\{\mathbb{P}_{v}:v_{r}=1\}$, respectively. It follows from Assouad’s lemma [1] that
$\sup_{\mathbb{P}\in\mathcal{P}}\mathbb{E}\lVert\widehat{f}^{[t]}-f^{[t]}\rVert^{2}_{\mathcal{L}^{2}}$ $\geq\sup_{\mathbb{P}\in\mathcal{P}_{0}}\mathbb{E}\lVert\widehat{f}^{[t]}-f^{[t]}\rVert^{2}_{\mathcal{L}^{2}}$ (12)
$\geq\frac{1}{2}C^{2}L_{m}^{2}(2m_{t})^{-2\alpha_{m}-1}\lVert\phi\rVert_{\mathcal{L}^{2}}^{2}\sum_{r=1}^{2m_{t}}\Big{(}1-\lVert\overline{\mathbb{P}}_{0,r}-\overline{\mathbb{P}}_{1,r}\rVert_{\mathrm{TV}}\Big{)}$
Notice that two mixtures ¯$\overline{\mathbb{P}}_{0,r}$ and ¯$\overline{\mathbb{P}}_{1,r}$ follow exactly the same distribution when none of the common design points $\{T_{j}^{[t]}:j=1,\ldots,m_{t}\}$ falls into the interval $[(r-1)/2m_{t},r/2m_{t})$. Since there are $m_{t}$ number of common design points, the Pigeonhole principle guarantees the number of such an $r\in\{1,\ldots,2m_{t}\}$ must be at least $m_{t}$. In other words, we have
$\sum_{r=1}^{2m_{t}}\Big{(}1-\lVert\overline{\mathbb{P}}_{0,r}-\overline{\mathbb{P}}_{1,r}\rVert_{\mathrm{TV}}\Big{)}\geq m_{t}.$
The desired Equation (6.1) is now immediate by combining Equation (12) and (6.1). ∎
6.2 Proof of Theorem 2.2
We have already shown in Theorem 2.1 that
$\sup_{\mathbb{P}\in\mathcal{P}}\mathbb{E}\lVert\widehat{f}_{\mathrm{CL}}^{[t]}-f^{[t]}\rVert^{2}_{\mathcal{L}^{2}}\lesssim L_{m}^{2}m_{t}^{-2\alpha_{m}}+\frac{\log^{2}n_{t}}{n_{t}}$
Thanks to the nature of $\widehat{f}^{[t]}$, it suffices to argue the following rate of the estimator $\widehat{f}_{\mathrm{TL}}^{[t]}$,
@@
Thus, it is sufficient to demonstrate the following propositions, with detailed proof available in the Supplementary material [5].
@@
It is now straightforward to choose the bandwidths $b_{s}=\Theta(m_{s}^{-1})$ and $b_{\delta}=\Theta(m_{t}^{-1})$ such that they satisfy $m_{s}b_{s}\geq 2C_{s}(d_{s}+1)$ and $m_{t}b_{\delta}\geq 2C_{t}(d_{\delta}+1)$, and to select any constant degrees $d_{s}\geq\omega(\alpha_{m})$ and $d_{\delta}\geq\omega(\alpha_{\delta})$. This concludes the proof. ∎
@@
We construct another sub-model $\mathcal{P}_{2}\subset\mathcal{P}$ as follows. The source samples $\mathcal{D}^{[s,1]},\ldots,\mathcal{D}^{[s,K]}$ are the most advantageous in the sense that $f^{[t]}=f^{[s,1]}=\cdots=f^{[s,K]}$. In other words, every sample $\mathcal{D}^{[t]},\mathcal{D}^{[s,1]},\ldots,\mathcal{D}^{[s,K]}$ is assumed to share the same data-generating process whose further specifications would be different depending on each case. As we have done in the proof of Theorem 2.1, let us take any $C^{\infty}$ function $\phi:\mathbb{R}\to\mathbb{R}$ such that $\phi^{(q)}(x)=0$ for any $q=0,1,2,\ldots$ and $x\not\in(0,1)$. We will prove Proposition 6.4 and 6.5 here. The proof of Proposition 6.6 and 6.7 will be presented separately in the Supplementary material [5]. ∎
@@
Under the sub-model $\mathcal{P}_{2}\subset\mathcal{P}$, let us assume the target random function $X^{[t]}_{1}$ as well as the source ones $X^{[s,1]}_{1},\ldots,X^{[s,K]}_{1}$ are unknown constant functions. Since their mean functions should coincide, the problem is fundamentally equivalent to estimating the scalar mean from $n_{t}+Kn_{s}$ number of independent and identically distributed observations under mean squared error. It is thus immediate to get the lower bound of the parametric rate:
@@
where $C>0$ is a small enough generic constant such that $g_{v}\in\mathcal{H}_{\alpha_{m}}(L_{m},M_{m})$. The sub-model $\mathcal{P}_{2}\subset\mathcal{P}$ consists of any probability measure $\mathbb{P}_{v}$ $(v\in\{0,1\}^{2m})$ whose common mean function $f^{[t]}=f^{[s,1]}$ is given as $g_{v}$. This sub-model $\mathcal{P}_{2}$ is Hamming separated in the sense that
@@
$\geq\frac{1}{2}C^{2}L_{m}^{2}(2m)^{-2\alpha_{m}-1}\lVert\phi\rVert_{\mathcal{L}^{2}}^{2}\sum_{r=1}^{2m}\Big{(}1-\lVert\overline{\mathbb{P}}_{0,r}-\overline{\mathbb{P}}_{1,r}\rVert_{\mathrm{TV}}\Big{)}$
Notice that two mixtures ¯$\overline{\mathbb{P}}_{0,r}$ and ¯$\overline{\mathbb{P}}_{1,r}$ follow exactly the same distribution when the interval $[(r-1)/2m,r/2m)$ contains nothing among $\{T_{j}^{[t]}:j=1,\ldots,m_{t}\}\cup\{T_{j}^{[s,1]}:j=1,\ldots,m_{s}\}$, the common design points for target and source domains. Since these design points are at most $m=m_{t}+m_{s}$, the number of such an $r\in\{1,\ldots,2m\}$ must be at least $m$ by Pigeonhole principle. In other words, we have

### v2

@@ 6 Proofs
Thanks to Theorem 2.1 and the nature of our estimator $\widehat{f}^{[t]}$, it suffices to argue the following rate of the transfer learning estimator $\widehat{f}_{\mathrm{TL}}^{[t]}$,
@@
We shall present the rates of convergence for those estimators, $\widehat{f}^{[s]}$ and $\widehat{\delta}^{[s]}$, in the following propositions. The proofs are given in the Supplementary material [5].
@@
It is therefore optimal to choose the bandwidths $b_{s}=\Theta(m_{s}^{-1})$ and $b_{\delta}=\Theta(m_{t}^{-1})$ such that they satisfy $m_{s}b_{s}\geq 2C_{s}(d_{s}+1)$ and $m_{t}b_{\delta}\geq 2C_{t}(d_{\delta}+1)$, and to select any constant degrees $d_{s}\geq\omega(\alpha_{m})$ and $d_{\delta}\geq\omega(\alpha_{\delta})$. This immediately concludes the proof. ∎
@@
We construct another sub-model $\mathcal{P}_{2}\subset\mathcal{P}$ as follows. The source samples $\mathcal{D}^{[s,1]},\ldots,\mathcal{D}^{[s,K]}$ are the most advantageous in the sense that $f^{[t]}=f^{[s,1]}=\cdots=f^{[s,K]}$. In other words, every sample $\mathcal{D}^{[t]},\mathcal{D}^{[s,1]},\ldots,\mathcal{D}^{[s,K]}$ is assumed to share the same data-generating process whose further specifications would be different depending on each case. Let us take any $C^{\infty}$ function $\phi:\mathbb{R}\to\mathbb{R}$ such that $\phi^{(q)}(x)=0$ for any $q=0,1,2,\ldots$ and $x\not\in(0,1)$. We will prove Proposition 6.3 and 6.4. The proof of Proposition 6.5 and 6.6 will be presented separately in the Supplementary material [5]. ∎
@@
Under the sub-model $\mathcal{P}_{2}\subset\mathcal{P}$, let us assume the target random function $X^{[t]}_{1}$ as well as the source ones $X^{[s,1]}_{1},\ldots,X^{[s,K]}_{1}$ are unknown constant functions. Since their mean functions should coincide, the problem is equivalent to estimating the scalar mean from $n_{t}+Kn_{s}$ number of independent and identically distributed observations under mean squared error. It is thus immediate to get the lower bound of the parametric rate:
@@
where $C>0$ is a small enough generic constant such that $g_{v}\in\mathcal{H}_{\alpha_{m}}(L_{m},M_{m})$. The sub-model $\mathcal{P}_{2}\subset\mathcal{P}$ consists of any probability measure $\mathbb{P}_{v}$ $(v\in\{0,1\}^{2m})$ whose common mean function $f^{[t]}=f^{[s,1]}$ is given as $g_{v}$. This sub-model $\mathcal{P}_{2}$ is Hamming separated in that
@@
$\geq\frac{1}{2}C^{2}L_{m}^{2}(2m)^{-2\alpha_{m}-1}\lVert\phi\rVert_{\mathcal{L}^{2}}^{2}\sum_{r=1}^{2m}\Big{(}1-\lVert\overline{\mathbb{P}}_{0,r}-\overline{\mathbb{P}}_{1,r}\rVert_{\mathrm{TV}}\Big{)},$
where $\lVert\,\cdot\,\rVert_{\mathrm{TV}}$ denotes the total variation distance. Notice that two mixtures ¯$\overline{\mathbb{P}}_{0,r}$ and ¯$\overline{\mathbb{P}}_{1,r}$ follow exactly the same distribution when the interval $[(r-1)/2m,r/2m)$ contains nothing among $\{T_{j}^{[t]}:j=1,\ldots,m_{t}\}\cup\{T_{j}^{[s,1]}:j=1,\ldots,m_{s}\}$, the common design points for target and source domains. Since these design points are at most $m=m_{t}+m_{s}$, the number of such an $r\in\{1,\ldots,2m\}$ must be at least $m$ by Pigeonhole principle. In other words, we have